# Statistical Taylor Expansion111

### v2

@@ 2 Statistical Taylor Expansion
2.1 The Uncorrelated Uncertainty Assumption [1]
[…]
figure Effect of noise on bit values of a measured value. The triangular wave signal and the added white noise are shown at top using the thin black line and the grey area, respectively. The values are measured by a theoretical 4-bit Digital-to-Analog Converter in ideal condition, assuming LSB is the 0th bit. The measured 3rd and 2nd bits without the added noise are shown using thin black lines, while the mean values of the measured 3rd and 2nd bits with the added noise are shown using thin grey lines. This figure has been published previously [1].
[…]
figure $\gamma$ vs $P$ for different $\gamma_{P}$ in Formula (2.4). If $\gamma_{P}$ is the maximal allowed correlation on the uncertainty level for uncorrelated uncertainty assumption to hold, then $\gamma$ is the maximal allowed correlation for two input to hold. For a given $\gamma_{P}$, $\gamma$ increases with finer $P$. This figure is a reproduction of Figure 2 in [1], with the P-axis in linear scale instead of in logarithmic scale, and with all texts improved.
@@ 2.1 The Uncorrelated Uncertainty Assumption [1]
However, the uncorrelated uncertainty assumption allows any two inputs to be highly correlated if each input is independently measured [1]:
Let $X$, $Y$, and $Z$ denote three mutually independent random variables [5] with variance $V(X)$, $V(Y)$ and $V(Z)$, respectively. Let $\alpha$ denote a constant. Let $Cov()$ denote the covariance function. Let $\gamma$ denote the correlation between $(X+Y)$ and $(\alpha X+Z)$. And let:
$\eta_{1}^{2}\equiv\frac{V(Y)}{V(X)};\;\;\;\eta_{2}^{2}\equiv\frac{V(Z)}{V(\alpha X)}=\frac{V(Z)}{\alpha^{2}V(X)};$ (2.1)
$\gamma=\frac{Cov(X+Y,\alpha X+Z)}{\sqrt{V(X+Y)}\sqrt{V(\alpha X+Z)}}=\frac{\alpha/|\alpha|}{\sqrt{1+\eta_{1}^{2}}\sqrt{1+\eta_{2}^{2}}}\equiv\frac{\alpha/|\alpha|}{1+\eta^{2}};$ (2.2)
Formula (2.2) gives the correlation $\gamma$ between two random variables, each of which contains a completely uncorrelated part and a completely correlated part $X$, with $\eta$ being the average ratio between these two parts. Formula (2.2) can also be interpreted reversely: if two random variables are correlated by $\gamma$, each of them can be viewed hypothetically as containing a completely uncorrelated part and a completely correlated part, with $\eta$ being the average ratio between these two parts  [^4] The widely used common factor analysis is also based on this principle.. The correlated parts common to different measurements are regarded as signals, which can either be desired or unwanted.
One special application of Formula (2.2) is the correlation between a measured signal and its true signal, in which noise is the uncorrelated part between the two. Figure 2.1 shows the effect of noise on the most significant two bits of a 4-bit measured signal when $\eta=1/4$. Its top chart shows a triangular waveform between 0 and 16 as a black line, and a white noise between -2 and +2, using the grey area. The measured signal is the sum of the triangle waveform and the noise. The middle chart of Figure 2.1 shows the values of the 3rd digit of the true signal as a black line, and the mean values of the 3rd bit of the measurement as a grey line. The 3rd bit is affected by the noise during its transition between 0 and 1. For example, when the signal is slightly below 8, only a small positive noise can turn the 3rd digit from 0 to 1. The bottom chart of Figure 2.1 shows the values of the 2nd digit of the signal and the measurement as a black line and a grey line, respectively. Figure 2.1 clearly shows that the correlation between the measurement and the true signal is less at the 2nd digit than at the 3rd digit. Quantitatively:
  • The overall measurement is 99.2% correlated to the signal with $\eta=1/8$;
  • The 3rd digit of the measurement is 97.0% correlated to the signal with $\eta=1/4$;
  • The 2nd digit of the measurement is 89.4% correlated to the signal with $\eta=1/2$;
  • The 1st digit of the measurement is 70.7% correlated to the signal with $\eta=1$;
  • The 0th digit of the measurement is 44.7% correlated to the signal with $\eta=2$.
The above conclusion agrees with the common experiences that, below the noise level of measured signals, noises rather than true signals dominate each digit.
Similarly, while the correlated portion between two values has the same value at each bit of the two values, the ratio of the uncorrelated portion to the correlated portion increases by 2-fold for each bit down from MSB of the two values. Quantitatively, let $P$ denote the precision of an imprecise value, and let $\eta_{P}$ denote the ratio of the uncorrelated portion to the correlated portion at level of uncertainty; then $\eta_{P}$ increases with decreased $P$ according to Formula (2.3). According to Formula (2.2), if two significant values are overall correlated with $\gamma$, at the level of uncertainty the correlation between the two values decreases to $\gamma_{P}$ according to Formula (2.4).
$\eta_{P}=\frac{\eta}{P},\;\;\;P<1;$ (2.3)
$\frac{1}{\gamma_{P}}-1=\left(\frac{1}{\gamma}-1\right)\frac{1}{P^{2}},\;\;\;P<1;$ (2.4)
Figure 2.1 plots the relation of $\gamma$ vs. $P$ for each given $\gamma_{P}$ in Formula (2.4). When $\gamma_{P}$ is less than a predefined maximal threshold (e.g., 2%, 5% or 10%), the two values can be deemed uncorrelated to each other at the level of uncertainty. For each independence standard $\gamma_{P}$, there is a maximal allowed correlation between two values below which the uncorrelated uncertainty assumption of statistical Taylor expansion holds. The maximal allowed correlation is a function of the larger precision of the two values according to Formula (2.4). Figure 2.1 shows that for two precisely measured values, their correlation $\gamma$ is allowed to be quite high.
In Formula (2.4), according to Formula (2.2), $P=\sqrt{P_{1}P_{2}}$, in which $P_{1}$ and $P_{2}$ are the precision of the two inputs, respectively. When $P=0$, $\gamma=1$, showing the implied statistical allowance when either or both inputs are precise.
@@ 2.3 Statistical Taylor Expansion
$=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\left(\zeta(n)-\zeta(j)\zeta(n-j)\right);$ (2.10)
When $\sigma$ is limited, Formula (2.7) provides statistical bounding $\tilde{x}\in(x-\sigma\delta x,x+\sigma\delta x)$.
@@
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\left(\zeta_{x}(m)\zeta_{y}(n)-\zeta_{x}(i)\zeta_{x}(m-i)\;\zeta_{y}(j)\zeta_{y}(n-j)\right);$ (2.13)
@@ 2.6 Dependency Tracing
  • Formula (2.30) shows $\delta^{2}(f+g)$, whose dependence tracing can be demonstrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.15). Formula (2.34) shows that Formula (2.3) uses Formula (2.30) between any two terms in the Taylor expansion of Formula (2.8). In contrast, Formula (2.35) shows that statistical Taylor expansion can no longer be treated as polynomial due to correlation between $x$ and $x^{n}$.
    $\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=$ $(\delta x)^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+(\delta x)^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$ (2.34)
@@ 3 Variance Arithmetic
3.1 Gaussian Input Uncertainty Assumption
The central limit theorem converges other distributions toward Gaussian [5] after addition and subtraction, and such convergence is fast [1]. In a digital computer, multiplication is implemented as a series of shift and addition, division is implemented as a series of shift and subtraction, and general function is calculated as addition of expansion terms [8][9]. This is why in general an uncertainty without bounds is assumed to be Gaussian distributed [2][3][5]. Variance arithmetic also assumes that an imprecise input for statistical Taylor expansion is Gaussian distributed.
In variance arithmetic, $\zeta(2n+1)=0$.
3.2 Bounding Factor and Bounding Leakage
$=$ $(2n-1)\zeta(\sigma,2n-2)-2N(\sigma)\sigma^{2n-1};$ (3.2)
The property of $\zeta(2n)$ when $\sigma^{2}\ll 2n$ is not sensitive to the input uncertainty distribution. Formula (3.4) is the $\zeta(2n)$ for uniform distribution between $[-\sigma,+\sigma]$. It is similar to Formula (3.3):
[…]
Formula (3.5) gives the probability $\epsilon$ of $\tilde{x}\not\in(-\sigma\delta x,+\sigma\delta x)$, in which $\xi(z)$ is the cumulative density function for Normal distribution [5]. $\epsilon$ is a probability defined as bounding leakage. When $\sigma=5$, $\epsilon=5.73\times 10^{-7}$, which is small enough for most applications. $\sigma=5$ is also a golden standard to assert a negative result [3].
$\epsilon=2-2\xi(\frac{1}{2}+\sigma);$ (3.5)
Thus, $\sigma=5$ in variance arithmetic.
@@ 3.4 Convergence, Stability, Validity, and Reliability
$\lim_{n\rightarrow+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}=P(x)^{2n}\sigma^{2n}\frac{\log(2n)}{(2n)^{2}}\sigma N(\sigma);$ (3.16)
@@ 4.2 Types of Uncertainties
The modelling errors arise when an approximate analytic solution is used, or when a real problem is simplified to obtain the solution. For example, Section 8 demonstrates that the discrete Fourier transformation is only an approximation for the mathematically defined continuous Fourier transformation. Conceptually, modeling errors originate from mathematics, so they are outside the domain for variance arithmetic.
@@ 8.3 Testing Signals
To avoid unfaithfulness of discrete Fourier transformation to continuous transformation, only Formula (8.1) and (8.2) with integer $n$ and $k$ will be used.

### v3

@@ 2 Statistical Taylor Expansion
2.1 The Uncorrelated Uncertainty Assumption
@@
If two signals are independently measured so that their measurement noises are not correlated. Their measurement precision are $P_{1}$ and $P2$, respectively. Even these two signal has overall correlation $\gamma$, at the level of the average precision $P=\sqrt{P_{1}P_{2}}$, the correlation is reduced to $\gamma_{P}$ according to Formula (2.1) [1].
$\frac{1}{\gamma_{P}}-1=\left(\frac{1}{\gamma}-1\right)\frac{1}{P^{2}};$ (2.1)
$\gamma_{P}$ decreases fast with finer $P$ so that when $P$ is fine enough, at the level of uncertainty, the two signals are independent of each other [1].
@@ 2.3 Statistical Taylor Expansion
$=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big{(}\zeta(n)-\zeta(j)\zeta(n-j)\big{)};$ (2.7)
When $\sigma$ is limited, Formula (2.4) provides statistical bounding $\tilde{x}\in(x-\sigma\delta x,x+\sigma\delta x)$. The probability $\epsilon$ for $\tilde{x}\not\in(x-\sigma\delta x,x+\sigma\delta x)$ is defined as bounding leakage.
@@
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\big{(}\zeta_{x}(m)\zeta_{y}(n)-\zeta_{x}(i)\zeta_{x}(m-i)\;\zeta_{y}(j)\zeta_{y}(n-j)\big{)};$ (2.10)
@@ 2.4 Bounding Symmetry and Asymptote
Empirically, as Formula (2.15), (2.18), and (2.23) will demonstrate, when $n\rightarrow+\infty$, $\zeta(n)\rightarrow\sigma^{n}/n$.
2.4.1 Uniform Input Uncertainty
For uniform distribution, the bounding factor is $\sigma=\sqrt{3}$, and the bounding leakage $\epsilon=0$. According to Formula (2.15), $\zeta(2n)=(\sqrt{3})^{2n}/(2n+1)$ while $\zeta(2n+1)=0$.
[…]
2.4.2 Gaussian Input Uncertainty
The central limit theorem converges other distributions toward Gaussian [5] after addition and subtraction, and such convergence is fast [1]. In a digital computer, multiplication is implemented as a series of shifts and additions, division is implemented as a series of shifts and subtractions, and general function is calculated as addition of expansion terms [8][9]. This is why in general an uncertainty without bounds is assumed to be Gaussian distributed [2][3][5].
For Gaussian uncertainty input, $\zeta(2n+1)=0$.
$=$ $(2n-1)\zeta(\sigma,2n-2)-2N(\sigma)\sigma^{2n-1};$ (2.17)
Formula (2.19) gives the bounding leakage $\epsilon$, in which $\xi(z)$ is the cumulative density function for Normal distribution [5].
$\epsilon=2-2\xi(\frac{1}{2}+\sigma);$ (2.19)
2.4.3 An Input Uncertainty with Limited Range
Formula (2.20) shows a distribution density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $[\sigma_{-}\lambda,\sigma\lambda]$ satisfying $0<\sigma_{-}<1<\sigma$. Formula (2.21) shows its mean-preserving equation. Formula (2.22) shows its variance momentum. Formula (2.23) shows its asymptote. Formula (2.24) shows its bounding leakage.
$\rho(\tilde{x},x,\delta x)$ $=\frac{\tilde{x}}{\lambda^{2}}e^{-\frac{\tilde{x}}{\lambda}};\;\;\;x=2\lambda;\;\;\;\delta^{2}x=2\lambda^{2};$ (2.20)
$e^{-(\sigma_{-})}(\sigma_{-})^{2}=e^{-\sigma}\sigma^{2};$ (2.21)
$\zeta(n)$ $=(n+1)!\left(e^{-(\sigma_{-})}\sum_{j=0}^{n+1}\frac{(\sigma_{-})^{j}}{j!}-e^{-\sigma}\sum_{j=0}^{n+1}\frac{\sigma^{j}}{j!}\right);$ (2.22)
$\lim_{n\rightarrow+\infty}\zeta(n)=\frac{\sigma^{n+2}}{n+2}e^{-\sigma};$ (2.23)
$\epsilon$ $=1-e^{-(\sigma_{-})}+e^{-\sigma};$ (2.24)
@@ 2.6 Dependency Tracing
  • Formula (2.37) shows $\delta^{2}(f+g)$, whose dependence tracing can be demonstrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.12). Formula (2.6) shows that Formula (2.3) uses Formula (2.37) between any two terms in the Taylor expansion of Formula (2.5).
    $\zeta(2n+1)=0:$
    $\;\;\;\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=(\delta x)^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+(\delta x)^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$
@@ 3 Variance Arithmetic
3.1 Choices of Bounding Factors
For discussion simplicity, only Gaussian input uncertainty will be considered in this paper.
For Gaussian input uncertainty, when $\sigma=5$, $\epsilon=5.73\times 10^{-7}$ according to Formula (2.19), which is small enough for most applications. $\sigma=5$ is also a golden standard to assert a negative result [3]. Thus, $\sigma=5$ in variance arithmetic for Gaussian input uncertainty.
@@ 3.3 Convergence, Stability, Validity, and Reliability
$\lim_{n\rightarrow+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}=P(x)^{2n}\sigma^{2n}\frac{2\log(2n)}{(2n)^{2}}\sigma N(\sigma);$ (3.11)
@@ 4.2 Types of Uncertainties
The modelling errors arise when an approximate analytic solution is used, or when a real problem is simplified to obtain the solution. For example, the predecessor of variance arithmetic [1] demonstrates that the discrete Fourier transformation is only an approximation for the mathematically defined continuous Fourier transformation. Conceptually, modeling errors originate from mathematics, so they are outside the domain for variance arithmetic.
@@ 8.3 Testing Signals
To avoid unfaithfulness of discrete Fourier transformation to continuous transformation [1], only Formula (8.1) and (8.2) with integer $n$ and $k$ will be used.